OpenStudy (anonymous):

Can someone please find the mass for each time? Time (sec) Mass (kg) Force (Newtons) 11000 8 9520 10 12000 7 11500 9

OpenStudy (anonymous):

force (newtons) = mass * acceleration f = m*a = (kg * m) / s^2